0,2,2,3,3,5,8,4,10,______what will be in the blank. Basically it is a mat problem supporting number series. I wanna know how can I solve this

Grade:9

2 Answers

Arun
6 years ago
Dear Avinandan
 
the missing no is 15
We have 2nd,5th ,8th no in sequence i.e 2,3,4
similarly 3rd , 6th ,9th no. are 2, 2+3=5, 5+5=10,
Similarly 1st,4th,7th and 10th no.0,0+3=3,3+5=8, 8+7=15
 
Regards
Arun (AskIITians Forum Expert)
Om
4 years ago
Devide the series in two parts like 2,3,4,5... And second part like 0,2,3,5,8,10,?,17 now 1^2-1,1^2+1,2^2-1,2^2+1,so and is 4^2-1